OpenStudy (anonymous):

A kite is flying at the end of a straight string that has a length of 250meters. The string makes an angle of 65degrees with the ground. How high above the ground is the kite?

OpenStudy (anonymous):



OpenStudy (anonymous):

right?

OpenStudy (anonymous):

That's what I drew too... Then do I use cos65?

OpenStudy (anonymous):

nope remember SOH CAH TOA?

OpenStudy (anonymous):

sine =opp/hyp cos =adj/hyp tan= opp/adj

OpenStudy (anonymous):

What do I use the? Sorry

OpenStudy (anonymous):

sine

OpenStudy (anonymous):

I thought that the side we were finding was the adjacent so I would use 250cos65=adjacent side

OpenStudy (anonymous):

adj would be the base side, h would be the opposite

OpenStudy (anonymous):

Thanks so much for your help!

OpenStudy (anonymous):

np gl